OpenStudy (gabylovesyou):

What is the logarithmic form of the equation e^3x ≈ 3247? log3x3247 = e ln 3247 = 3x 3 logxe = 3247 ln 3x = 3247

satellite73 (satellite73):

log base 3 is written\(\ln(x)\)

jimthompson5910 (jim_thompson5910):

I think @satellite73 meant to say `log base e is written ln(x)`

OpenStudy (gabylovesyou):

oh so its D

OpenStudy (gabylovesyou):

right ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let's find out Apply natural log to both sides \[\Large e^{3x} \approx 3247\] \[\Large \ln\left(e^{3x}\right) \approx \ln\left(3247\right)\] \[\Large 3x*\ln\left(e\right) \approx \ln\left(3247\right)\] \[\Large 3x*1 \approx \ln\left(3247\right)\] \[\Large 3x \approx \ln\left(3247\right)\] \[\Large \ln\left(3247\right) \approx 3x\] I used the idea that `ln(e) = 1` and the idea that `ln(x^y) = y*ln(x)`

OpenStudy (gabylovesyou):

@jim_thompson5910 Thank youu